Wardell Company purchased a mainframe on January 1, 2016, at a cost of $53,000. The computer was depreciated using the straight-line method over an estimated five-year life with an estimated residual value of $5,000. On January 1, 2018, the estimate of useful life was changed to a total of 10 years, and the estimate of residual value was changed to $800.

Required:
1. Prepare the year-end journal entry for depreciation in 2018. No depreciation was recorded during the year.
2. Prepare the year-end journal entry for depreciation in 2018. Assume that the company uses the sum-of-the-years' -digits method instead of the straight-line method.
Answer #1

Solution:

1)Journal entry:

Original depreciation =(53,000 - 5000/5) = 9600 per year

Accumlated depreciation = 900*2 = 19,200

Revised depreciation =(53,000 - 19,200-800)/8 = 4,125

Date Accounts & explanation Debit Credit
2018 Depreciation expenses 4,125
Accumlated depreciation 4,125
(To record depreciation)

2)

Journal entry:

Sum of year digit = 5+4+3+2+1 = 15

Accumlated depreciation(2016 to 2017) =(53,000 - 5,000)*9/15 = 28,800

Revised sum of year digit = 8+7+6+5+4+3+2+1 = 36

2018 depreciation =(53,000 - 28,800 - 800)*8/36 = 5,200

Date Accounts & explanation Debit Credit
2018 Depreciation expense 5,200
Accumlated depreciation 5,200
(To record depreciation)